Three weights are placed on a plank. One of 15 kg. mass is placed 0.6 meters in from one end, the next of 12 kg. mass is placed 1.5 meters in from the same end and the last of 18 kg. mass is placed 3 meters from this end. If the mass of the plank be ignored, find the resultant moment about the end of the plank
1
Expert's answer
2020-09-24T11:08:32-0400

The resultant moment is the algebraic sum of all moments:


"\tau=F_1d_1+F_2d_2+F_3d_3,\\\space\\
\tau=g(m_1d_1+m_2d_2+m_3d_3)=\\
=9.8(15\cdot0.6+12\cdot1.5+18\cdot3)=794\text{ N}\cdot\text{m}"